OpenStudy (anonymous):

WILL AWARD MEDAL! Find if possible, the absolute maximum and minimum values of the given function on the indicated interval. f(x)=sin(x+pi/2) on (0, 7pi/4) Steps please!

OpenStudy (anonymous):

hmm.... is the (0,7pi/4) a co-ordinate?

OpenStudy (anonymous):

does it help to know that \[\sin(x+\frac{\pi}{2})=\cos(x)\]?

OpenStudy (anonymous):

you know you have to be a bit careful here

OpenStudy (anonymous):

if it is the open interval \((0,\frac{7\pi}{4})\) the function \(\cos(x)\) actually has no maximum value

OpenStudy (anonymous):

since it approaches 1 at \(x=0\) but \(0\) is not in the interval so the correct answer would be "minimum value is \(-1\), but no maximum value"

OpenStudy (anonymous):

Hi can you help me with the steps @satellite73 and the answer that was given with the problem is maximum value f(0)=1; minimum value of f(pi)=-1